# Re(l)ality:

### v1

@@ Abstract
Using the fiber bundle framework, this work investigates the conceptual and mathematical foundations of reference frames in General Relativity by contrasting two paradigms. The View from Nowhere interprets frame representations as perspectives on an invariant equivalence class, while the View from Everywhere posits each frame representation as constituting reality itself. This conception of reality is termed ”Relality.” The paper critically examines the philosophical and practical implications of these views, with a focus on reconciling theory with experimental practice. Central to the discussion is the challenge of providing a perspicuous characterisation of ontology. The View from Nowhere aligns with the so-called ‘sophisticated approach on symmetries’ and it complicates the empirical grounding of theoretical constructs. In contrast, the View from Everywhere offers a relational ontology that avoids the abstraction of equivalence classes.
@@ 1 Introduction
In a diffeomorphism-invariant theory, variables can be broadly categorized into two types: (1) highly non-local quantities, which are defined across the entire spacetime, such as integrals over global regions, and (2) relational quantities, which are constructed by correlating field values. This work primarily focuses on the second category, commonly referred to as relational observables.
@@
The first, the, View from Nowhere is the dominant perspective in the literature. It asserts that changing a reference frame does not affect the physical content of the description, as a reference frame provides merely a perspective on a shared, objective reality. This idea has its roots in Special Relativity, where different inertial reference frames offer distinct perspectives on a Lorentz-invariant physical reality. Prominent examples of this view include the perspective-neutral framework introduced by Vanrietvelde et al. (2020) and related work by Giacomini et al. (2019) and Kabel et al. (2024).
@@
Once reference frames are introduced, it is straightforward to construct relational quantities. Analogously as done above, we can distinguish various possible formulations of relational objects in GR, when written in terms of reference frames. Here, I provisionally denote a reference frame without further specification as a set of 4 scalar quantities $\phi^{(I)}_{I=1,\dots,4}$ associating a real quadruple with each point $P$ of the manifold. (See the next section for a more detailed discussion). We can distinguish between relational quantities that are:
@@ 2. II)
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality typical of the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’).
@@ 2 The Bundle Formalism: A Gauge Perspective on Reference Frames
Choosing a reference frame can also be interpreted as selecting a specific gauge. In fact, the choice of a gauge is equivalent to the choice of a reference frame (see Dittrich (2007) for insights on the relationship between gauge-fixed observables and relational observables). In gauge theories with a principal fiber bundle structure, gauge-fixing determines a section through the fiber bundle. Given a symmetry group $S$, each fiber corresponds to a gauge orbit — the set of all configurations of a field that are related by gauge transformations, generated by the constraints of the theory. Specifically, there exists a one-to-one correspondence between each gauge orbit and the equivalence class of models under the symmetry $s\in S$. More concretely, for $m\in M$ and $s\in S$, the orbit $\mathcal{O}_{m}$ is defined as $\mathcal{O}_{m}=\{m_{s}\mid s\in S\}$, where $m_{s}$ denotes the model to which the transformation $s$ is applied. Because of the theory’s symmetry, the space of models includes significant redundancy. The physical content of the theory is captured by a single representative from each gauge orbit, while other models within the same orbit are redundant, isomorphic copies. [^3] This does not imply that gauge freedom constitutes mere ”descriptive fluff” (see Earman (2004)). On the contrary, the correspondence between gauge-fixing and reference frame selection highlights the relational nature of physics. The additional degrees of freedom are meaningful as they represent the possible ways a system can form observables relative to another system. Isomorphic models provide ”handles” through which systems can couple; see Rovelli (2014) and Adlam (2024). This redundancy is typically resolved by fixing a gauge. Similarly, in GR, one fixes a reference frame through a coordinate gauge condition (see Bamonti (2023); Gomes (2024b)).
@@
This encapsulates the concept of a ‘relational, gauge-invariant observable’ and a ‘reference frame.’ [^8] Strictly, a ‘relational observable’ is not automatically gauge-invariant (Bamonti and Gomes, 2024b).
@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
If $g_{ab}$ satisfies the condition described in Equation (3) for some $F_{\phi_{r}}$ (resp. $F_{\phi_{b}}$), then any pair $(g_{ab},\phi_{r}^{(I)})$ (resp. $(g_{ab},\phi_{b}^{(I)})$), can be mapped to a unique reference frame representation of the metric. This mapping is achieved through the action of $f_{\phi_{r}}$ (resp. $f_{\phi_{b}}$), yielding $(g_{ab})_{F_{\phi_{r}}}:=f_{\phi_{r}}^{*}g_{ab}$, which is equivalent to the local gauge-invariant observable $[\phi_{r}^{-1}]^{*}g_{ab}:=g_{IJ}(\phi_{r})$. The same is for for $\phi_{b}$. Refer to Figure 2.
@@
As per Equation (2), $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ are gauge-invariant, relational observables that represent distinct physical scenarios. Nonetheless, they are frequently interpreted as mere perspectives on a shared equivalence class $[g_{ab}]$ of isometric metrics. This equivalence class is regarded as the sole structure with ‘ontological significance’, meaning it is considered the fundamental structure underlying physical reality. As shown in Table 1, $[g_{ab}]$ is frame-free, spatiotemporally implicit, and gauge-invariant. This interpretation constitutes what I call:
@@
Moreover, Dewar argues that, Sophistication, unlike Eliminativism, does not require an intrinsic characterization of $[M]$ for an ontological commitment to equivalence classes alone. However, Martens and Read (2020) have criticised this method of defining invariant structures — defined as the interpretational approach by Møller-Nielsen (2017) — for failing to ensure a perspicuous understanding of the shared ontology underlying various models.
@@
I have already expressed many of my qualms about the View from Nowhere, and in the next section I will offer an alternative on how to interpret the formalism of relational observables. Before doing so, however, I want to dwell on a further problem related to the View from Nowhere: namely the loss of physical information. It can be said that, within this view, different choices of reference frames can be conceptualised as ‘windows of knowledge that provide partial views of a shared invariant structure’, available to an observer (see Adlam (2024) on how to schematise a conscious observer in a diff-invariant theory such as GR). Accordingly, an observer can adjust her perspective to explore specific aspects of this invariant structure, effectively selecting one reference frame over another based on her investigative focus.
However, this approach implies a loss of the complete, ’absolute’ information contained within the ontologically fundamental structure, which is defined by the equivalence class. This idea resonates with Einstein (1917)’s assertion that “ […]a definite choice of the system of reference […] is contrary to the spirit of the relativity principle.” Similarly, (Adlam, 2024, p.9) argues that “diffeomorphism invariance could finally be broken by the observer herself.” Geng (2024) also echoes this viewpoint.
@@ 4 The View from Everywhere: Frame representations are all that exist
Clearly, this represents a passive diffeomorphism transformation, demonstrating how changes in reference frames preserve the underlying physical quantities while altering their representation. [^15] Due to the one-to-one correspondence between active and passive diffeomorphisms, from the active perspective, the relationship between $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ can be interpreted as the external diffeomorphism $\textbf{d}:=\phi_{r}^{-1}\circ\textbf{m}\circ\phi_{r}$, or equivalently $\textbf{d}:=\phi_{b}\circ\textbf{m}^{-1}\circ\phi_{b}^{-1}$. In GR, local gauge-invariant observables are covariant under frame transformations (see Pitts (2022)). [^16] Pitts actually refers to covariance relative to a coordinate system. Here, we extend this notion to encompass reference frames.
@@
What conception of reality emerges from this framework? Within the View from Everywhere, each of the frame-dependent physical situations constitutes what I call Relality. This concept emphasises that reality is fundamentally local and relational. The concept of Relality is intriguing and deserves to be analysed further.
Drawing on the Kantian perspective used to interpret the View from Nowhere, I argue that the View from Everywhere fundamentally rejects the notion of a shared, implicit noumenal reality underlying its phenomenological manifestations. According to the View from Everywhere, there is no ’hidden’ structure beyond these frame-specific representations; each gauge-invariant quantity constitutes the entirety of what fundamentally exists. Accordingly, the two observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ do not describe two distinct phenomenal perspectives of a shared, ontologically independent noumenal reality represented by $[g_{ab}]$. Instead, each constitutes a Relality and there is nothing else.
@@
As already stressed, the Sophistication approach posits that the fundamental description of physical reality should rely on equivalence classes. The View from Everywhere proves that this assumption is unnecessary. Gauge-invariant relational observables, such as $g(\phi_{r})$ and $g(\phi_{b})$, represent distinct, fully-fledged physical scenarios, leaving no basis to hypothesise the existence of an additional ’hidden’ structure beyond them. This interpretation aligns with a principle of ”ontological parsimony”, which I sustain. The View from Everywhere, being ontologically more conservative, posits no existence beyond the Relalities expressible within reference frames.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices181818m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
@@
In contrast, in the context of the View from Everywhere, a choice of reference frame constitutes an empirical choice, rather than a conventional one. This is because we cannot conventionally choose another reference frame that represents the ‘same’, ‘true’ physical situation. A different choice is in all respects a different Relality.

### v2

@@ Abstract
Using the fiber bundle framework, this work investigates the conceptual and mathematical foundations of reference frames in General Relativity by contrasting two paradigms. The View from Nowhere interprets frame representations as perspectives on an invariant equivalence class, while the View from Everywhere posits each frame representation as constituting a fully-fledged reality itself. What emerges is a conception of reality that I term ”Relality.” The paper critically examines the philosophical and practical implications of these views, with a focus on reconciling theory with experimental practice. Central to the discussion is the challenge of providing a perspicuous characterisation of ontology. The View from Nowhere aligns with the so-called ‘ sophisticated approach to symmetries’ and complicates the empirical grounding of theoretical constructs. In contrast, the View from Everywhere offers a relational ontology that avoids the abstraction of equivalence classes.
@@ 1 Introduction
In a diffeomorphism-invariant theory, variables can be broadly categorized into two types: (1) highly non-local quantities, which are defined across the entire spacetime, such as integrals over global regions, and (2) relational quantities, which are constructed by correlating field values. This work focusses primarily on the second category, commonly referred to as relational observables.
@@
The first, the, View from Nowhere is the dominant perspective in the literature. It asserts that changing a reference frame does not affect the physical content of the description, as a reference frame provides merely a perspective on a shared, objective reality. This idea has its roots in Special Relativity, where different inertial reference frames offer distinct perspectives on a Lorentz-invariant physical reality. Prominent examples of this view include the perspective-neutral framework introduced by Vanrietvelde et al. (2020) and related work by Giacomini et al. (2019) and Kabel et al. (2024). Importantly, I will use the term ‘perspective’ in the ordinary deflated sense throughout the paper. [^2] The term ’perspective’ is widely used in the philosophy of science. For example, Michela Massimi uses the notion of perspective in terms of a historically and culturally situated epistemic lens within a debate on scientific realism Massimi (2022).
@@
Once reference frames are introduced, it is straightforward to construct relational quantities. Analogously as done above, we can distinguish various possible formulations of relational objects in GR, when written in terms of reference frames. Here, I provisionally denote a reference frame without further specification as a set of 4 scalar quantities $\phi^{(I)}\mid_{I=1,\dots,4}$ associating a real quadruple with each point $P$ of the manifold. (See the next section for a more detailed discussion). We can distinguish between relational quantities that are:
@@ 2. II)
NB: of course, a frame explicit object will be also spatiotemporally explicit, in the specific sense enclosed by the relational locality typical of the use of reference frames. To avoid confusion, I have decided to distinguish between frame-explicit and spatiotemporally explicit objects, preserving the term ‘spatiotemporal’ to indicate objects written in terms of manifold points (if it helps, the reader may also interpret a spatiotemporally explicit object as ‘manifold explicit’). [^4] The concept of spacetime lacks a universally agreed definition, and its interpretation varies across different frameworks. It may be understood as: (i) the manifold $\mathcal{M}$; (ii) the combination $(\mathcal{M},g_{ab})$, which consists of the manifold equipped with the metric field representing the gravitational field; or (iii) the gravitational field $g_{ab}$ alone. In interpretations (i) and (ii), the distinction lies in whether $\mathcal{M}$ is considered an ontologically independent entity—a stage upon which dynamical variables act—or whether it is inseparable from the fields. Interpretation (iii), however, treats $\mathcal{M}$ as a purely mathematical construct without ontological status (see Rovelli and Gaul (2000); Rovelli (2006); Rovelli and Vidotto (2015); Einstein et al. (2015)). An additional perspective considers spacetime as an emergent structure, defined in terms of observables in the Dirac (gauge-invariant) sense. These observables correspond to the happenings of events, which specify the ”when” and ”where” of physical phenomena. In this paper I choose the more conservative option and identify space-time with option (ii).
@@ 2 The Bundle Formalism: A Gauge Perspective on Reference Frames
Choosing a reference frame can also be interpreted as selecting a specific gauge. In fact, the choice of a gauge is equivalent to the choice of a reference frame (see Dittrich (2007) for insights on the relationship between gauge-fixed observables and relational observables). In gauge theories with a principal fiber bundle structure, gauge-fixing determines a section through the fiber bundle. Given a symmetry group $S$, each fiber corresponds to a gauge orbit — the set of all configurations of a field that are related by gauge transformations, generated by the constraints of the theory. Specifically, there exists a one-to-one correspondence between each gauge orbit and the equivalence class of models under the symmetry $s\in S$. More concretely, for $m\in M$ and $s\in S$, the orbit $\mathcal{O}_{m}$ is defined as $\mathcal{O}_{m}=\{m_{s}\mid s\in S\}$, where $m_{s}$ denotes the model to which the transformation $s$ is applied. Due to the symmetry of the theory, the model space includes significant redundancy. The physical content of the theory is captured by a single representative from each gauge orbit, while other models within the same orbit are redundant, isomorphic copies. [^5] This does not imply that gauge freedom constitutes mere ”descriptive fluff” (see Earman (2004)). On the contrary, the correspondence between gauge-fixing and reference frame selection highlights the relational nature of physics. The additional degrees of freedom are meaningful as they represent the possible ways a system can form observables relative to another system. Isomorphic models provide ”handles” through which systems can couple; see Rovelli (2014) and Adlam (2024). This redundancy is typically resolved by fixing a gauge. Similarly, in GR, one fixes a reference frame through a coordinate gauge condition (see Bamonti (2023); Gomes (2024b)).
@@
This encapsulates the concept of a ’relational, gauge-invariant observable’ and a ’reference frame’. [^10] Strictly, a ‘relational observable’ is not automatically gauge-invariant (Bamonti and Gomes, 2024b).
@@ 3 The View from Nowhere: Frame representations are perspectives on an equivalence class
If $g_{ab}$ satisfies the condition described in Equation (3) for some $F_{\phi_{r}}$ (resp. $F_{\phi_{b}}$), then any pair $(g_{ab},\phi_{r}^{(I)})$ (resp. $(g_{ab},\phi_{b}^{(I)})$), can be mapped into a unique reference frame representation of the metric. This mapping is achieved through the action of $f_{\phi_{r}}$ (resp. $f_{\phi_{b}}$), yielding $(g_{ab})_{F_{\phi_{r}}}:=f_{\phi_{r}}^{*}g_{ab}$, which is equivalent to the local gauge-invariant observable $[\phi_{r}^{-1}]^{*}g_{ab}:=g_{IJ}(\phi_{r})$. The same is for for $\phi_{b}$. Refer to Figure 2.
@@
As per Equation (2), $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ are gauge-invariant, relational observables that represent distinct physical scenarios. Nonetheless, they are frequently interpreted as mere perspectives on a shared equivalence class $[g_{ab}]$ of isometric metrics. This equivalence class is regarded as the sole structure with ‘ontological significance’, which means that it is considered the fundamental structure underlying physical reality. As shown in Table 1, $[g_{ab}]$ is frame-free, spatiotemporally implicit, and gauge-invariant. This interpretation constitutes what I call:
@@
Moreover, Dewar argues that, Sophistication, unlike Eliminativism, does not require an intrinsic characterisation of $[M]$ for an ontological commitment to equivalence classes alone. However, Martens and Read (2020) have criticised this method of defining invariant structures — defined as the interpretational approach by Møller-Nielsen (2017) — for failing to ensure a perspicuous understanding of the shared ontology underlying various models.
@@
I have already expressed many of my qualms about the View from Nowhere, and in the next section I will offer an alternative on how to interpret the formalism of relational observables. Before doing so, however, I want to dwell on a further problem related to the View from Nowhere: namely the loss of physical information. [^17] This loss is exemplified by what Tong says about the choice of a gauge: The [gauge] redundancy allows us to make manifest the properties of quantum field theories, such as unitarity, locality, and Lorentz invariance, that we feel are vital for any fundamental theory of physics but which teeter on the verge of incompatibility. If we try to remove the redundancy by fixing some specific gauge, some of these properties will be brought into focus, while others will retreat into murk. By retaining the redundancy, we can flit between descriptions as is our want, keeping whichever property we most cherish in clear sight. (Tong, 2018, p.1) It can be said that, within this view, different choices of reference frames can be conceptualised as ‘windows of knowledge that provide partial views of a shared invariant structure’, available to an observer (see Adlam (2024) on how to schematise a conscious observer in a diff-invariant theory such as GR). Accordingly, an observer can adjust her perspective to explore specific aspects of this invariant structure, effectively selecting one reference frame over another based on her investigative focus.
However, this approach implies a loss of the complete, ’absolute’ information contained within the ontologically fundamental structure, which is defined by the equivalence class.
This idea also resonates with Einstein (1917)’s assertion that “ […]a definite choice of the system of reference […] is contrary to the spirit of the relativity principle.” Similarly, (Adlam, 2024, p.9) argues that “diffeomorphism invariance could finally be broken by the observer herself.” Geng (2024) also echoes this viewpoint.
@@ 4 The View from Everywhere: Frame representations are all that exist
Clearly, this represents a passive diffeomorphism transformation, demonstrating how changes in reference frames preserve the underlying physical quantities while altering their representation. [^18] Due to the one-to-one correspondence between active and passive diffeomorphisms, from the active perspective, the relationship between $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ can be interpreted as the external diffeomorphism $\textbf{d}:=\phi_{r}^{-1}\circ\textbf{m}\circ\phi_{r}$, or equivalently $\textbf{d}:=\phi_{b}\circ\textbf{m}^{-1}\circ\phi_{b}^{-1}$. In GR, local gauge-invariant observables are covariant under frame transformations (see Pitts (2022)). [^19] Pitts actually refers to covariance relative to a coordinate system. Here, I extend this notion to encompass reference frames. Note that this does not mean that there are no frame-invariant quantities in general, only that these quantities are not local (see Table 1).
@@
What conception of reality emerges from this framework? Within the View from Everywhere, each of the frame-dependent physical situations constitutes a fully-fledged reality. The collection of such realities constitutes what I call Relality. The concept of Relality is intriguing and deserves to be analysed further. Importabtly, this concept helps me to emphasise that reality is fundamentally local and relational.
Drawing on the Kantian perspective used to interpret the View from Nowhere, I argue that the View from Everywhere fundamentally rejects the notion of a shared, implicit noumenal reality underlying its phenomenological manifestations. According to the View from Everywhere, there is no ’hidden’ structure beyond these frame-specific representations; each gauge-invariant quantity constitutes the entirety of what fundamentally exists. Accordingly, the two observables $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$ do not describe two distinct phenomenal perspectives of a shared, ontologically independent noumenal reality represented by $[g_{ab}]$. Instead, each constitutes a reality and there is nothing else. This is because the concept of Relality, formalised by $[g_{IJ}]$, does not represent a shared ontology. Its utility lies only in ensuring the inter-translatability of each reality. This supports the abandonment of the concept of ‘perspective’.
@@
As already stressed, the Sophistication approach posits that the fundamental description of physical reality should rely on equivalence classes. The View from Everywhere proves that this assumption is unnecessary. Gauge-invariant relational observables, such as $g_{IJ}(\phi_{r})$ and $g_{IJ}(\phi_{b})$, represent distinct, fully-fledged physical scenarios, leaving no basis to hypothesise the existence of an additional ’hidden’ structure beyond them. This interpretation aligns with a principle of ”ontological parsimony”, which I sustain. The View from Everywhere, being ontologically more conservative, posits no ‘entities’ beyond the realities expressible within reference frames.
@@
First, the breaking of the diffeomorphic freedom of theory is precluded by the existence of the external diffeomorphism m. This map relates different reference frame choices212121m serves as the analogue of the transition map between sections of a principal bundle, given by $\mathrm{\bf{t}}_{\sigma\sigma^{\prime}}(\varphi)=f_{\sigma^{\prime}}(\varphi)^{-1}f_{\sigma}(\varphi)$ (see Gomes (2024a))., ensuring GR to remain m-covariant (see also Bamonti and Gomes (2024a)).
@@
In contrast, in the context of the View from Everywhere, a choice of reference frame constitutes an empirical choice, rather than a conventional one. This is because we cannot conventionally choose another reference frame that represents the ‘same’, ‘true’ physical situation. A different choice is in all respects a different reality.